4M 21.) A satellite A moves around a planet of mass4M, another satellite B moves around a planet of m the orbital period of satellite A in terms of that of satellite B if the orbital radius is the same for B.1-12TB E. TA-4T
Answer #1

Orbital Period is given by

T=2\pi \sqrt{\frac{a^{3}}{GM}}

=>\frac{T_{A}}{T_{B}}=\sqrt{\frac{M_{B}}{M_{A}}}=\sqrt{\frac{M}{4M}}=\frac{1}{2}

TA=(1/2)TB